# OmniDocBench: Benchmarking Diverse PDF Document Parsing with Comprehensive Annotations

### v1

@@ 4 OmniDocBench Evaluation Methodology
To provide a fair and comprehensive evaluation for various models, we proposed an end-to-end evaluation pipeline consisting of several modules, including extraction, matching algorithm, and metric calculation, as shown in Figure 3. It ensures that OmniDocBench automatically performs unified evaluation on end-to-end DCE tasks, thereby producing reliable and effective evaluation results.
@@ 4.1 Extraction
Preprocessing: The model-generated markdown text should be preprocessed, which includes removing images, eliminating markdown tags at the beginning of the document, and standardizing the number of repeated characters.
Special Component Extraction: Extraction is primarily carried out using regular expression matching. To ensure that the extraction of content does not interfere with each other, it is necessary to follow a specific order. The extraction sequence is as follows: LaTeX tables, HTML tables, display formulas, markdown tables (which are then converted into HTML format), and code blocks.
Pure Text Extraction: After extracting special components, the remaining content is considered pure text. Paragraphs are separated by double line breaks, allowing them to participate in subsequent matching processes, thus aligning with reading order annotation units in the GTs. If no double line break exists, single line breaks are used for paragraph separation. Additionally, previously extracted code blocks are merged into the text category for processing.
Inline Formula Format Converting: We standardized inline formulas within paragraphs to Unicode format. This was necessary because different models produce inconsistent outputs for inline formulas. For formulas originally written in Unicode, it is hard to extract them using regular expressions. Therefore, to ensure a fair comparison, we do not extract inline formulas for separate evaluation. Instead, we include them in their Unicode format alongside the text paragraphs for evaluation.
Reading Order Extraction: Upon completion of the extraction, the start and end positions of the extracted content in the original markdown are recorded for subsequent reading order calculation.
@@
Pipeline Tools MinerU 0.058 0.211 0.278 0.577 66.9 49.5 79.4 62.7 0.305 0.461 0.079 0.288 0.180 0.384
Marker 0.141 0.303 0.667 0.868 18.4 12.7 54.0 45.8 0.718 0.763 0.138 0.306 0.416 0.560
Mathpix 0.101 0.358 0.306 0.454 71.4 72.7 77.9 68.2 0.322 0.416 0.105 0.275 0.209 0.376
Expert VLMs GOT-OCR 0.187 0.315 0.360 0.528 81.8 51.4 53.5 48.0 0.521 0.594 0.141 0.28 0.302 0.429
Nougat 0.365 0.998 0.488 0.941 17.4 16.9 40.3 0.0 0.622 1.000 0.382 0.954 0.464 0.973
General VLMs GPT4o 0.144 0.409 0.425 0.606 76.4 48.2 72.8 63.7 0.363 0.474 0.128 0.251 0.265 0.435
Qwen2-VL 0.252 0.251 0.468 0.572 54.9 60.9 59.9 66.8 0.591 0.587 0.255 0.223 0.392 0.408
InternVL2 0.353 0.290 0.543 0.701 69.8 49.6 63.8 61.1 0.616 0.638 0.317 0.228 0.457 0.464
@@
Model Type Models Book Slides Financial Report Textbook Exam Paper Magazine Academic Papers Notes Newspaper Average
Pipeline Tools MinerU 0.044 0.124 0.033 0.102 0.159 0.072 0.025 0.984 0.148 0.188
Marker 0.188 0.327 0.087 0.292 0.423 0.134 0.102 0.470 0.270 0.255
Mathpix 0.131 0.168 0.202 0.199 0.278 0.138 0.091 0.631 0.648 0.276
Expert VLMs GOT-OCR 0.105 0.222 0.067 0.132 0.204 0.198 0.179 0.388 0.771 0.252
Nougat 0.734 0.958 1.000 0.820 0.930 0.83 0.214 0.991 0.871 0.816
General VLMs GPT4o 0.157 0.163 0.348 0.187 0.281 0.173 0.146 0.607 0.751 0.313
Qwen2-VL 0.094 0.08 0.145 0.148 0.219 0.065 0.315 0.298 0.79 0.239
InternVL2 0.216 0.098 0.162 0.184 0.247 0.150 0.419 0.226 0.903 0.289
@@
Models Fuzzy Water Color Mean Variance
Pipeline Tools
MinerU 0.15 0.151 0.107 0.136 0.0004
Marker 0.286 0.436 0.290 0.337 0.0049
Mathpix 0.294 0.290 0.182 0.255 0.0027
Expert VLMs
GOT-OCR 0.175 0.190 0.186 0.184 0.0000
Nougat 0.934 0.915 0.873 0.907 0.0006
General VLMs
GPT4o 0.263 0.195 0.184 0.214 0.0012
Qwen2-VL 0.101 0.157 0.114 0.124 0.0006
InternVL2 0.120 0.197 0.155 0.157 0.0010
Table 4: End-to-end text recognition on OmniDocBench: evaluation under various page attributes using the edit distance metric. Columns represent: Fuzzy (Fuzzy scan), Water (Watermark), Color (Colorful background).
Models Single Double Three Complex Mean Variance
Pipeline Tools
MinerU 0.311 0.101 0.117 0.376 0.226 0.0143
Marker 0.231 0.251 0.309 0.378 0.292 0.0033
Mathpix 0.189 0.175 0.225 0.413 0.250 0.0091
Expert VLMs
GOT-OCR 0.163 0.145 0.257 0.468 0.258 0.0165
Nougat 0.852 0.601 0.662 0.873 0.747 0.0139
General VLMs
GPT4o 0.109 0.204 0.254 0.426 0.248 0.0132
Qwen2-VL 0.098 0.248 0.517 0.429 0.323 0.0263
InternVL2 0.082 0.312 0.682 0.444 0.380 0.0472
Table 5: End-to-end reading order evaluation on OmniDocBench: results across different column layout types using Normalized Edit Distance.
@@ 4.2 Matching Algorithm
To avoid the impact of paragraph splitting on the final results, we proposed a method, Adjacency Search Match, that merges and splits paragraphs in both GTs and Preds to achieve the best possible match. The specific strategy involves: i) Calculate a metrix of Normalized Edit Distance between GTs and Preds. If the similarity between a Pred and a GT exceeds a specific threshold, they are considered a successful match. ii) For the rest, we apply fuzzy matching to determine whether one string is a subset of another string. If so, we further apply the truncation and merging algorithm which would try to merge adjacent paragraph. This process would continue to merge more paragraph until the Normalized Edit Distance starts to decrease. After this process, the best match will be found for GTs and Preds.
4.3 Metric Calculation
Ignore Handling: We implement an ignore logic for certain components in PDF page content, meaning they participate in matching but are excluded from metric calculations. This is mainly because of inconsistent output standards among models, which should not affect the validation results. For fairness, we ignore: (1) Headers, footers, page numbers, and page footnotes, which are handled inconsistently by different models. (2) Captions for figures, tables, and footnotes often have uncertain placement, complicating reading order. Additionally, some models embed table captions in HTML or LaTeX tables, while others treat them as plain text.
Metric: Different calculation methods are used for various document components: (1) Pure Text: We calculate Normalized Edit Distance, averaging these metrics at the sample level to obtain the final scores. (2) Tables: All tables are converted to HTML format before calculating the TEDS metric and Normalized Edit Distance. (3) Formulas: Formulas are currently evaluated using the CDM [35], Normalized Edit Distance, and BLEU. We did not convert interline formulas into Unicode because Unicode cannot represent certain complex formulas, such as matrices. (4) Reading Order: Reading order use the Normalized Edit Distance as metric. It only involves text components, where tables, images, and ignored components do not participate in the final reading order calculation.
5 Benchmarks
Model Book Slides Research Report Textbook Exam Paper Magazine Academic Literature Notes Newspaper Average mAP
DiT-L 43.44 13.72 45.85 15.45 3.40 29.23 66.13 0.21 23.65 26.90
LayoutLMv3 42.12 13.63 43.22 21.00 5.48 31.81 64.66 0.80 30.84 28.84
DOCX-Chain 30.86 11.71 39.62 19.23 10.67 23.00 41.60 1.80 16.96 21.27
DocLayout-YOLO 43.71 48.71 72.83 42.67 35.40 51.44 66.84 9.54 57.54 48.71
@@ 5.1 Component-specific Evaluation Results
OCR-based Models PaddleOCR 76.8 71.8 80.1 67.9 74.3 81.1 74.5 70.6/75.2 71.3/74.1 72.7/74.0 23.3/74.6 73.6
RapidTable 80.0 83.2 91.2 83.0 79.7 83.4 78.4 77.1/85.4 76.7/83.9 77.6/84.9 25.2/83.7 82.5
Expert VLMs StructEqTable 72.0 72.6 81.7 68.8 64.3 80.7 85.0 65.1/76.8 69.4/73.5 66.8/75.7 44.1/73.3 72.7
GOT-OCR 72.2 75.5 85.4 73.1 72.7 78.2 75.7 65.0/80.2 64.3/77.3 70.8/76.9 8.5/76.3 74.9
General VLMs Qwen2-VL-7B 70.2 70.7 82.4 70.2 62.8 74.5 80.3 60.8/76.5 63.8/72.6 71.4/70.8 20.0/72.1 71.0
InternVL2-8B 70.9 71.5 77.4 69.5 69.2 74.8 75.8 58.7/78.4 62.4/73.6 68.2/73.1 20.4/72.6 71.5
@@
Expert Vision Models PaddleOCR 0.071 0.055 0.118 0.060 0.038 0.085 0.060 0.015 0.285 0.021
Tesseract OCR 0.179 0.553 0.553 0.453 0.463 0.394 0.448 0.369 0.979 0.982
Surya 0.057 0.123 0.164 0.093 0.186 0.235 0.104 0.634 0.767 0.255
GOT-OCR 0.041 0.112 0.135 0.092 0.052 0.155 0.091 0.562 0.966 0.097
Mathpix 0.033 0.240 0.261 0.185 0.121 0.166 0.180 0.038 0.185 0.638
Vision Language Models Qwen2-VL 0.072 0.274 0.286 0.234 0.155 0.148 0.223 0.273 0.721 0.067
InternVL2 0.074 0.155 0.242 0.113 0.352 0.269 0.132 0.610 0.907 0.595
GPT4o 0.020 0.224 0.125 0.167 0.140 0.220 0.168 0.115 0.718 0.132
@@
GOT-OCR 74.1 28.0 55.07 0.290
Mathpix 86.6 2.8 66.56 0.322
Pix2Tex 73.9 39.5 46.00 0.337
UniMERNet-B 85.0 60.2 60.84 0.238
GPT4o 86.8 65.5 45.17 0.282
InternVL2 67.4 54.5 47.63 0.308
Qwen2-VL 83.8 55.4 53.71 0.285
@@
The OmniDocBench dataset features comprehensive and precise annotations, allowing for a fair and rigorous comparison of various document content extraction algorithms in real-world scenarios. Based on the distinct characteristics of these algorithms, we categorize document content extraction methods into three main classes:
Pipeline Tools. These methods integrate layout detection and various content recognition tasks (such as OCR, table recognition, and formula recognition) into a document parsing pipeline for content extraction. Prominent examples include MinerU [36], Marker [30], and Mathpix  [^4] https://mathpix.com/.
Expert VLMs. These are large multimodal models specifically trained for document parsing tasks. Representative models include GOT-OCR2.0 [39] and Nougat [5].
General VLMs. These are general-purpose large multimodal models inherently capable of document parsing. Leading models in this category include GPT-4o [2], Qwen2-VL [38], and InternVL2 [6].
@@ 5.2 End-to-End Evaluation Results
Utilizing the OmniDocBench dataset and our evaluation framework, we conducted end-to-end assessments of mainstream document parsing methods, evaluating their performance from input PDF images to the resultant document parsing outputs.
Overall Evaluation Results. As illustrated in Table 2, pipeline tools specifically designed for document parsing, demonstrate superior performance across the board. MinerU and Mathpix achieved the best results for English and Chinese pages, respectively. In contrast, even the best general-purpose Vision Language Models (VLMs), GPT-4o, exhibits a performance gap compared to these specialized models, especially in Chinese. This trend is evident across sub-tasks like text recognition, formula recognition, and table recognition, where methods tailored for document parsing consistently outperform others. This advantage is largely due to the fine-tuning of these models on large datasets specific to document parsing tasks.
@@
Performance Across Diverse Page Types. To gain deeper insights into model performance on diverse document types, we evaluated text recognition tasks across different page types. As shown in Table 3, an intriguing finding emerged: For commonly used data, such as academic papers and financial reports, pipeline tools perform well. However, for more specialized data like slides and handwritten notes, general VLMs demonstrate stronger generalization. The reason is clear: Pipeline tools and expert VLMs are relatively more constrained by the range of training data, whereas general VLMs having been trained on a wide variety of samples, maintained excellent recognition performance even in traditionally challenging long-tail scenarios, underscoring the value of VLMs.
Performance on Pages with Specific Attributes. For documents in OmniDocBench with attributes such as fuzzy scans, watermarks, and colorful backgrounds, our evaluation results are presented in Table 4. In these scenarios, the VLMs InternVL2 and Qwen2-VL exhibit the strongest resistance to interference, achieving the best accuracy and robustness. MinerU also performs commendably.
Performance on Different Column Layout Types. OmniDocBench annotates page attributes such as column layout type, which is crucial for analyzing model performance in reading order. As depicted in Table 5, all models experience a noticeable decline in reading order accuracy when dealing with complex layouts. MinerU and Mathpix excels in reading order across various column layouts, demonstrating robust performance across different page types.
From these end-to-end evaluations, it is evident that pipeline tools like MinerU and Mathpix, specifically designed for document parsing, achieve the best overall performance. However, in terms of versatility and scalability, VLMs offer a distinct advantage over pipeline tools. Fine-tuning a general large model like Qwen2-VL with specialized data could yield models even more adept at document parsing, indicating a promising direction for future research in multimodal approaches.
The OmniDocBench dataset provides comprehensive annotations for document parsing, including layout detection, text boxes and content, formula boxes and content, and table boxes and content. These detailed annotations enable the evaluation of current state-of-the-art (SOTA) methods across various document types, allowing us to analyze their performance in diverse scenarios. Additionally, these results can be used to assemble enhanced pipeline tools for document parsing tasks.
5.3 Single Algorithm Evaluation Results

### v2

@@ 4 OmniDocBench Evaluation Methodology
To provide a fair and comprehensive evaluation for various models, we proposed an end-to-end evaluation pipeline consisting of several modules, including extraction, matching algorithm, and metric calculation, as shown in Figure 4. It ensures that OmniDocBench automatically performs unified evaluation on document parsing, thereby producing reliable and effective evaluation results.
@@ 4.1 Extraction
Preprocessing. The model-generated markdown text should be preprocessed, which includes removing images, eliminating markdown tags at the beginning of the document, and standardizing the number of repeated characters.
Elements Extraction. Extraction is primarily carried out using regular expression matching. To ensure that the extraction of elements does not interfere with each other, it is necessary to follow a specific order. The extraction sequence is as follows: LaTeX tables, HTML tables, display formulas, markdown tables (which are then converted into HTML format), and code blocks.
Pure Text Extraction. After extracting special components, the remaining content is considered pure text. Paragraphs are separated by double line breaks, allowing them to participate in subsequent matching processes, thus aligning with reading order annotation units in the GTs. If no double line break exists, single line breaks are used for paragraph separation. Additionally, previously extracted code blocks are merged into the text category for processing.
Inline Formula Format Converting. We standardized inline formulas within paragraphs to Unicode format. This was necessary because different models produce inconsistent outputs for inline formulas. For formulas originally written in Unicode, it is hard to extract them using regular expressions. Therefore, to ensure a fair comparison, we do not extract inline formulas for separate evaluation. Instead, we include them in their Unicode format alongside the text paragraphs for evaluation.
Reading Order Extraction. Upon completion of the extraction, the start and end positions of the extracted content in the original markdown are recorded for subsequent reading order calculation.
@@
Pipeline Tools MinerU [42] 0.061 0.215 0.278 0.577 57.3 42.9 78.6 62.1 0.18 0.344 0.079 0.292 0.15 0.357
Marker [34] 0.08 0.315 0.53 0.883 17.6 11.7 67.6 49.2 0.619 0.685 0.114 0.34 0.336 0.556
Mathpix  [^4] https://mathpix.com/ 0.105 0.384 0.306 0.454 62.7 62.1 77.0 67.1 0.243 0.32 0.108 0.304 0.191 0.365
Expert VLMs GOT-OCR [45] 0.189 0.315 0.360 0.528 74.3 45.3 53.2 47.2 0.459 0.52 0.141 0.28 0.287 0.411
Nougat [7] 0.365 0.998 0.488 0.941 15.1 16.8 39.9 0.0 0.572 1.000 0.382 0.954 0.452 0.973
General VLMs GPT4o [2] 0.144 0.409 0.425 0.606 72.8 42.8 72.0 62.9 0.234 0.329 0.128 0.251 0.233 0.399
Qwen2-VL-72B [44] 0.096 0.218 0.404 0.487 82.2 61.2 76.8 76.4 0.387 0.408 0.119 0.193 0.252 0.327
InternVL2-76B [8] 0.353 0.290 0.543 0.701 67.4 44.1 63.0 60.2 0.547 0.555 0.317 0.228 0.44 0.443
@@
Model Type Models Book Slides Financial Report Textbook Exam Paper Magazine Academic Papers Notes Newspaper Overall
Pipeline Tools MinerU [42] 0.055 0.124 0.033 0.102 0.159 0.072 0.025 0.984 0.171 0.206
Marker [34] 0.074 0.34 0.089 0.319 0.452 0.153 0.059 0.651 0.192 0.274
Mathpix 4 0.131 0.22 0.202 0.216 0.278 0.147 0.091 0.634 0.69 0.3
Expert VLMs GOT-OCR [45] 0.111 0.222 0.067 0.132 0.204 0.198 0.179 0.388 0.771 0.267
Nougat [7] 0.734 0.958 1.000 0.820 0.930 0.83 0.214 0.991 0.871 0.806
General VLMs GPT4o [2] 0.157 0.163 0.348 0.187 0.281 0.173 0.146 0.607 0.751 0.316
Qwen2-VL-72B [44] 0.096 0.061 0.047 0.149 0.195 0.071 0.085 0.168 0.676 0.179
InternVL2-76B [8] 0.216 0.098 0.162 0.184 0.247 0.150 0.419 0.226 0.903 0.3
@@
Models Fuzzy Water Color None
MinerU [42] 0.15/0.048 0.151/0.031 0.107/0.052 0.079/0.035
Marker [34] 0.333/0.092 0.484/0.126 0.319/0.127 0.062/0.125
Mathpix 4 0.294/0.064 0.290/0.059 0.216/0.09 0.135/0.043
GOT-OCR [45] 0.175/0.05 0.190/0.056 0.186/0.097 0.177/0.081
Nougat [7] 0.934/0.051 0.915/0.071 0.873/0.096 0.615/0.208
GPT4o [2] 0.263/0.078 0.195/0.057 0.184/0.078 0.186/0.072
Qwen2-VL-72B [44] 0.082 /0.01 0.172/ 0.078 0.104/0.05 0.084/0.042
InternVL2-76B [8] 0.120/0.013 0.197/0.042 0.155/0.059 0.261/0.082
Table 4: End-to-end text recognition on OmniDocBench: evaluation under various page attributes using the edit distance metric. The value is Mean/Variance of scores in the attribute group. Columns represent: Fuzzy (Fuzzy scan), Water (Watermark), Color (Colorful background). None (No special issue)
Models Single Double Three Complex
MinerU [42] 0.311/0.187 0.101/0.013 0.117/0.046 0.385/0.057
Marker [34] 0.299/0.143 0.299/0.299 0.149/0.063 0.363/0.086
Mathpix 4 0.207/0.123 0.188/0.07 0.225/0.029 0.452/0.177
GOT-OCR [45] 0.163/0.106 0.145/0.059 0.257/0.072 0.468/0.185
Nougat [7] 0.852/0.084 0.601/0.224 0.662/0.093 0.873/0.09
GPT4o [2] 0.109/0.112 0.204/0.076 0.254/0.046 0.426/0.188
Qwen2-VL-72B [44] 0.066/0.048 0.145/0.049 0.204/0.055 0.394/0.203
InternVL2-76B [8] 0.082/0.052 0.312/0.069 0.682/0.098 0.444/0.174
Table 5: End-to-end reading order evaluation on OmniDocBench: results across different column layout types using Normalized Edit Distance. The value is Mean/Variance of scores in the attribute group.
@@ 4.2 Matching Algorithm
Adjacency Search Match. To avoid the impact of paragraph splitting on the final results, we proposed Adjacency Search Match, that merges and splits paragraphs in both GTs and Preds to achieve the best possible match. The specific strategy involves: i) Calculate a metrix of Normalized Edit Distance between GTs and Preds. The Pred and GT pairs whose similarity exceeds a specific threshold are considered as successful match. ii) For the rest, we apply fuzzy matching to determine whether one string is a subset of another string. If so, we further apply the merging algorithm which would try to merge adjacent paragraph. This process would continue to merge more paragraph until the Normalized Edit Distance starts to decrease. After this process, the best match will be found for GTs and Preds.
Ignore Handling. We implement an ignore logic for certain components in PDF page content, meaning they participate in matching but are excluded from metric calculations. This is mainly because of inconsistent output standards among models, which should not affect the validation results. For fairness, we ignore: (1) Headers, footers, page numbers, and page footnotes, which are handled inconsistently by different models. (2) Captions for figures, tables, and footnotes often have uncertain placements, thus complicating the reading order. Additionally, some models embed table captions in HTML or LaTeX tables, while others treat them as plain text.
Model Backbone Params Book Slides Research Report Textbook Exam Paper Magazine Academic Literature Notes Newspaper Average
DiT-L [24] ViT-L 361.6M 43.44 13.72 45.85 15.45 3.40 29.23 66.13 0.21 23.65 26.90
LayoutLMv3 [17] RoBERTa-B 138.4M 42.12 13.63 43.22 21.00 5.48 31.81 64.66 0.80 30.84 28.84
DocLayout-YOLO [53] v10m 19.6M 43.71 48.71 72.83 42.67 35.40 51.44 64.64 9.54 57.54 47.38
SwinDocSegmenter [4] Swin-L 223M 42.91 28.20 47.29 32.44 20.81 52.35 48.54 12.38 38.06 35.89
GraphKD [5] R101 44.5M 39.03 16.18 39.92 22.82 14.31 37.61 44.43 5.71 23.86 27.10
DOCX-Chain [50] - - 30.86 11.71 39.62 19.23 10.67 23.00 41.60 1.80 16.96 21.27
@@
OCR-based Models PaddleOCR[23] 76.8 71.8 80.1 67.9 74.3 81.1 74.5 70.6/75.2 71.3/74.1 72.7/74.0 23.3/74.6 73.6
RapidTable[37] 80.0 83.2 91.2 83.0 79.7 83.4 78.4 77.1/85.4 76.7/83.9 77.6/84.9 25.2/83.7 82.5
Expert VLMs StructEqTable[55] 72.8 75.9 83.4 72.9 76.2 76.9 88.0 64.5/81.0 69.2/76.6 72.8/76.4 30.5/76.2 75.8
GOT-OCR [45] 72.2 75.5 85.4 73.1 72.7 78.2 75.7 65.0/80.2 64.3/77.3 70.8/76.9 8.5/76.3 74.9
General VLMs Qwen2-VL-7B [44] 70.2 70.7 82.4 70.2 62.8 74.5 80.3 60.8/76.5 63.8/72.6 71.4/70.8 20.0/72.1 71.0
InternVL2-8B [8] 70.9 71.5 77.4 69.5 69.2 74.8 75.8 58.7/78.4 62.4/73.6 68.2/73.1 20.4/72.6 71.5
@@
Expert Vision Models PaddleOCR [23] 0.071 0.055 0.118 0.060 0.038 0.085 0.060 0.015 0.285 0.021
Tesseract OCR  [^5] https://github.com/tesseract-ocr/tesseract 0.179 0.553 0.553 0.453 0.463 0.394 0.448 0.369 0.979 0.982
Surya  [^6] https://github.com/VikParuchuri/surya 0.057 0.123 0.164 0.093 0.186 0.235 0.104 0.634 0.767 0.255
GOT-OCR [45] 0.041 0.112 0.135 0.092 0.052 0.155 0.091 0.562 0.966 0.097
Mathpix 4 0.033 0.240 0.261 0.185 0.121 0.166 0.180 0.038 0.185 0.638
Vision Language Models Qwen2-VL-72B [44] 0.072 0.274 0.286 0.234 0.155 0.148 0.223 0.273 0.721 0.067
InternVL2-76B [8] 0.074 0.155 0.242 0.113 0.352 0.269 0.132 0.610 0.907 0.595
GPT4o [2] 0.020 0.224 0.125 0.167 0.140 0.220 0.168 0.115 0.718 0.132
@@
GOT-OCR [45] 74.1 28.0 55.07 0.290
Mathpix 4 86.6 2.8 66.56 0.322
Pix2Tex  [^7] https://github.com/lukas-blecher/LaTeX-OCR 73.9 39.5 46.00 0.337
UniMERNet-B [40] 85.0 60.2 60.84 0.238
GPT4o [2] 86.8 65.5 45.17 0.282
InternVL2-76B [8] 67.4 54.5 47.63 0.308
Qwen2-VL-72B [44] 83.8 55.4 53.71 0.285
@@
4.3 Metric Calculation
Pure Text. We calculate Normalized Edit Distance [21], averaging these metrics at the sample level to obtain the final scores.
Tables. All tables are converted to HTML format before calculating the Tree-Edit-Distance-based Similarity (TEDS) [54] metric and Normalized Edit Distance.
Formulas. Formulas are currently evaluated using the Character Detection Matching (CDM) metric [41], Normalized Edit Distance, and BLEU [33].
Reading Order. Reading order is evaluated using the Normalized Edit Distance as metric. It only involves text components, with tables, images, and ignored components excluded from the final reading order calculation.
5 Benchmarks
Based on the distinct characteristics of these algorithms, we categorize document content extraction methods into three main classes:
  • Pipeline Tools: These methods integrate layout detection and various content recognition tasks (such as OCR, table recognition, and formula recognition) into a document parsing pipeline for content extraction. Prominent examples include MinerU [42] (v0.9.3), Marker [34] (v1.2.3), and Mathpix4.
  • Expert VLMs: These are large multimodal models specifically trained for document parsing tasks. Representative models include GOT-OCR2.0 [45] and Nougat [7].
  • General VLMs: These are general-purpose large multimodal models inherently capable of document parsing. Leading models in this category include GPT-4o [2], Qwen2-VL-72B [44], and InternVL2-76B [8].
@@ 5.1 End-to-End Evaluation Results
Overall Evaluation Results. As illustrated in Table 2, pipeline tools such as MinerU and Mathpix, demonstrate superior performance across sub-tasks like text recognition, formula recognition, and table recognition. Moreover, the general Vision Language Models (VLMs), Qwen2-VL, and GPT4o, also exhibit competitive performance. Almost all algorithms score higher on English than on Chinese pages.
@@
Performance Across Diverse Page Types. To gain deeper insights into model performance on diverse document types, we evaluated text recognition tasks across different page types. Intriguingly, as shown in Table 3, pipeline tools perform well for commonly used data, such as academic papers and financial reports. Meanwhile, for more specialized data, such as slides and handwritten notes, general VLMs demonstrate stronger generalization. Notably, most VLMs fail to recognize when dealing with the Newspapers, while pipeline tools achieve significantly better performance.
Performance on Pages with Visual Degradations. In Table 4, we further analyze performance on pages containing common document-specific challenges, including fuzzy scans, watermarks, and colorful backgrounds. VLMs like InternVL2 and Qwen2-VL exhibit higher robustness in these scenarios despite visual noise. Among pipeline tools, MinerU remains competitive due to its strong layout segmentation and preprocessing capabilities.
Performance on Different Layout Types. Page layout is a critical factor in document understanding, especially for tasks involving reading order. OmniDocBench annotates layout attributes such as single-column, multi-column, and complex custom formats. Across all models, we observe a clear drop in accuracy on multi-column and complex layouts. MinerU shows the most consistent reading order prediction, though its performance dips on handwritten single-column pages due to recognition noise.
Discussion on End-to-End Results. 1) While general VLMs often lag behind specialized pipelines and expert models on standard documents (e.g., academic papers), they generalize better to unconventional formats (e.g., notes) and perform more robustly under degraded conditions (e.g., fuzzy scans). This is largely due to their broader training data, enabling better handling of long-tail scenarios compared to models trained on narrow domains. 2) VLMs, however, struggle with high-density documents like newspapers due to limitations in input resolution and token length. In contrast, pipeline tools leverage layout-based segmentation to process components individually, maintaining accuracy in complex layouts. Enhancing VLMs with layout-aware designs and domain-specific fine-tuning offers a promising path forward. OmniDocBench facilitates this by providing detailed annotations for layout, text, formulas, and tables, enabling comprehensive benchmarking and modular tool development for diverse document parsing tasks.
5.2 Single Task Evaluation Results